### v1

# The Role of Graph Topology in the Performance of Biomedical Knowledge Graph Completion Models

Abstract
Knowledge Graph Completion has been increasingly adopted as a useful method for several tasks in biomedical research, like drug repurposing or drug-target identification. To that end, a variety of datasets and Knowledge Graph Embedding models has been proposed over the years. However, little is known about the properties that render a dataset useful for a given task and, even though theoretical properties of Knowledge Graph Embedding models are well understood, their practical utility in this field remains controversial. We conduct a comprehensive investigation into the topological properties of publicly available biomedical Knowledge Graphs and establish links to the accuracy observed in real-world applications. By releasing all model predictions and a new suite of analysis tools we invite the community to build upon our work and continue improving the understanding of these crucial applications.
Machine Learning, ICML, Knowledge Graphs, Knowledge Graph Completion, Graph Topology, Drug Discovery, Biomedical Research
1 Introduction
Knowledge Graphs (KGs) have become a powerful way of integrating information about a domain into a single structured representation. KGs represent knowledge as a collection of facts in the form of triples $(h,r,t)$, which make a statement about the relationship $r$ between two entities $h$ and $t$. In KGs from the biomedical domain, entities are genes, diseases, drugs, or pathways (among others), whilst relations represent biological interactions or associations between them (MacLean, 2021; Bonner et al., 2022a; Zeng et al., 2022). Biomedical KGs have been used to support a variety of tasks within the complex and costly drug discovery process, such as drug repurposing (Bang et al., 2023), identifying synergistic drug combinations (Zitnik et al., 2018) and target discovery (Paliwal et al., 2020). Unlike in other domains, biomedical KGs frequently comprise information from a wide range of conceptual abstraction levels. Edges representing information with a low abstraction level, such as experimentally validated protein-protein interactions (Protein A, interacts, Protein B), are found in the same KG as mid abstraction level edges, such as associations between genes and diseases taken from a genome-wide association study (Gene A, associates, Disease X), and high abstraction level edges, such as human-curated ontology hierarchies (Disease X, subclass of, Disease Y). This makes biomedical KGs uniquely interesting to consider when they are used in combination with Machine Learning algorithms, as models must learn to make predictions across these different abstraction levels simultaneously.
Since KGs, including biomedical KGs, are typically incomplete, inferring missing triples is a key Machine Learning application – a task known as Knowledge Graph Completion in the literature. Perhaps the most common method of making predictions with KGs involves using Knowledge Graph Embedding (KGE) models, which learn low dimensional representations for all entities and relation types in a KG and combine them through a scoring function to compute the likelihood of a missing edge. In the biomedical domain, Knowledge Graph Completion has been used for a diverse set of tasks including predicting new links between genes and diseases (Paliwal et al., 2020). However, little is known about the practical ability of ML models to leverage topological patterns in the graph in order to make better predictions. An example of such a pattern is composition, where a relation between two entities is inferred by the presence of a shared third entity. In the biomedical domain, this could be the case when a drug is inferred to treat a disease because they share a common gene connection. Ideally, we would like to discover new drug repurposing opportunities by learning from such patterns in the data. Similarly, the inference pattern uses the presence of a certain relation type between an entity pair to infer the presence of a different relation between the same pair, which could be used to discover novel therapeutic gene-disease connections by aggregating multiple evidence types.
While the theoretical implications of these topological patterns for KGE models are well understood – in fact, new KGE models are often designed to capture key topological patterns – their impact on the predictive performance in real-world Knowledge Graph Completion tasks remains an open question. Recent work has started to investigate this (Ali et al., 2021; Jin et al., 2023; Teneva & Hruschka, 2023), but has primarily focussed on the properties of relation types rather than individual triples. When comparing or proposing new KGE models, this approach helps to understand which topological patterns can be captured by the relation embedding. However, when investigating the impact of the graph structure on KGE model performance, this relation level approach might neglect critical micro-level structural data. Moreover, existing work has not focussed specifically on the biomedical domain, where there is a clear need for understanding the relationship between KG topology and model performance. New KGE models are typically developed and evaluated on a small set of general-domain KGs, which may contain distinct topological patterns compared to those from the biomedical domain. Biomedical KGs are often constructed by practitioners to address a specific purpose, requiring decisions about the key entities and relation types to include, the optimal data sources from which to extract information and how best to model the desired knowledge in a KG schema. This manual graph composition process is costly and time-consuming, an issue that can be compounded when KGs are designed as input for ML models, often requiring several iterations to best tune the graph to the predictive task at hand (Bonner et al., 2022a; Ratajczak et al., 2022). Thus, there is a growing interest in understanding how any decisions taken which affect topological properties of a KG can ultimately impact predictions made from them.
In this paper, we explore the relationship between biomedical KG topology and KGE model accuracy. Our main contributions are as follows:
  • We provide an in-depth analysis of the topological properties of six public KGs, focussing on the biomedical domain, and compare the corresponding predictive performance of four well-established KGE models. While previous work has limited the investigation at the macro-level of relation types, we are able to detect stronger patterns linking topological properties to predictive accuracy by zooming in at the level of individual triples.
  • We look in detail at highly relevant relations that biomedical practitioners are most interested in inferring. We give evidence of the topological differences in how they are represented in different KGs and how this reflects on predictive performance.
  • To aid future work in this area, we propose a standardized framework to describe KG topology and release a dedicated Python package kg-topology-toolbox implementing it [^1] https://github.com/graphcore-research/kg-topology-toolbox.
Finally, to help bridge the gap between industry and academy, we make the predictions of the trained models available to the community (Cattaneo et al., 2024) to conduct further analysis.
2 Related Work
Since the adoption of representation learning for Knowledge Graph Completion, there has been interest in linking predictive performance of models to topological structures in the underlying Knowledge Graphs (Bordes et al., 2013). These structures, taking into account the heterogeneous relation types in KGs, have been split into two categories: edge cardinalities and topological patterns (see Section 3 for the definitions used in this study). Indeed, it has become common for new models to be designed with the capacity of capturing these patterns (Cao et al., 2021; Chao et al., 2021; Sun et al., 2019), working under the assumption that better modelling them will result in improved predictive performance (Jin et al., 2023).
Edge cardinality patterns group relations based on the cardinalities of their head and tail entity sets. Four cardinality patterns have been proposed: one-to-one, one-to-many, many-to-one, and many-to-many (Bordes et al., 2013; Wang et al., 2014; Yang et al., 2015). On the other hand, topological patterns categorise relationships based on underlying logical principles that define how entities interact. Several such patterns have been proposed, with common examples being symmetric/anti-symmetric, inverse and composition (Sun et al., 2019; Ali et al., 2021). In the literature, it is common to consider these patterns at the level of whole relation types, meaning that all triples with a given relation type must conform to a certain pattern for it to be said to be present.
Whilst both pattern types are frequently discussed when comparing architectures, quantifying how model predictive performance correlates with them is less common. Existing work has typically been limited in scale, often only performed on a new approach when it is proposed. For example, the original TransE paper investigated the impact of cardinality patterns on model performance, finding that models could more easily predict heads in one-to-many relations and tails in many-to-one relations (Bordes et al., 2013). The DistMult model observed similar results, whilst also measuring the ability of the model to capture composition relations (Yang et al., 2015). TransH was found to be better at capturing symmetric relations (Wang et al., 2014). RotatE demonstrated better predictive performance than existing models on composition relation types, whilst the authors also found implicit evidence of other topological patterns in the embeddings (Sun et al., 2019).
There has been limited work providing a comprehensive study on the impact of these patterns across different models and datasets, although some studies have started to appear recently. For example, an investigation across 3 common benchmark KGs (FB15k-237, WN18RR and Yago310) and 21 KGE models found that symmetric relations were easier to predict that anti-symmetric and compositions ones (Ali et al., 2021). Another study compared the performance of 7 KGE models on FB15k-237 and WN18RR, focussing on 4 different topological patterns (Jin et al., 2023), finding that a model’s theoretical capacity to capture a certain pattern does not guarantee good predictive performance on it. However, none of these studies explores the impact of edge patterns on biomedical KGs specifically.
KGE models have been shown to be strongly impacted by entity degree, with models often being heavily biased towards highly connected entities (Bonner et al., 2022b; Rossi et al., 2021b, a; Mohamed et al., 2020). However, disentangling the effect of degree from other topological properties when considering model performance is often overlooked by prior studies. One study showed that as entity degree increases, performance on symmetric relations decreases (Jin et al., 2023). However, it is still an open research question as to how degree obscures other correlations in the performance of models on different topological patterns.
The prevalence at which different topological patterns occur in real-world datasets is also an active area of consideration. Teneva & Hruschka (2023) investigated the frequency of topological and cardinality patterns in 29 real-world KGs from a range of domains, including biomedical. They observed that, under their chosen pattern definitions, many relations in biomedical KGs are predominantly many-to-many and that there was a lack of composition relations. Other work investigated the prevalence of symmetric and composition relations in 3 common KG benchmark datasets, finding that the majority of relations were anti-symmetric and that composition relations were rarer (Ali et al., 2021).
It is also crucial to understand how the presence of topological patterns could impact validation metrics if not considered when creating train/test splits. The benchmark dataset FB15k was found to contain inverse triple pairs where one part was in the train set and its counterpart in the test set, meaning that models could simply memorise pairs (Toutanova & Chen, 2015). A new dataset, FB15k-237, was created to remove these train/test set leakages. Another common dataset, WN18, was found to have a similar problem, with the WN18RR dataset created to address this (Toutanova & Chen, 2015).
3 Knowledge Graph Topological Properties
For a triple $(h,r,t)$ in a KG $\mathcal{G}$ we define the head out-degree as the cardinality of the set $\{\hat{t}\mid\exists\hat{r}:(h,\hat{r},\hat{t})\in\mathcal{G}\}$ and, analogously, the tail in-degree as the cardinality of $\{\hat{h}\mid\exists\hat{r}:(\hat{h},\hat{r},t)\in\mathcal{G}\}$. We further define the head out-degree of same relation $\textrm{deg}_{r}(h)$ and the tail in-degree of same relation $\textrm{deg}_{r}(t)$ as the cardinalities of $\{\hat{t}\mid(h,r,\hat{t})\in\mathcal{G}\}$ and $\{\hat{h}\mid(\hat{h},r,t)\in\mathcal{G}\}$, respectively.
The edge cardinality of a triple $(h,r,t)$ is then defined as follows:
$\textrm{deg}_{r}(h)$ $\textrm{deg}_{r}(t)$ $=1$ $>1$
$=1$ one-to-one many-to-one
$>1$ one-to-many many-to-many
Examples of the four possible edge cardinalities are provided using a small sample graph in Figure 1.
Figure 1: Example of edge cardinalities in a KG with two relation types (blue, red).
We also define the following edge topological patterns for triples $(h,r,t)\in\mathcal{G}$:
  1. a)
    $(h,r,t)$ is symmetric $\iff h\neq t$ and $(t,r,h)\in\mathcal{G}$;
  2. b)
    $(h,r,t)$ has inference $\iff\exists r^{\prime}\neq r:(h,r^{\prime},t)\in\mathcal{G}$;
  3. c)
    $(h,r,t)$ has inverse $\iff\exists r^{\prime}\neq r:(t,r^{\prime},h)\in\mathcal{G}$;
  4. d)
    $(h,r,t)$ has composition $\iff\exists r_{1},r_{2}$ and $n\notin\{h,t\}:(h,r_{1},n),(n,r_{2},t)\in\mathcal{G}$.
[…]
(a) Symmetry
[…]
(b) Inference
[…]
(c) Inverse
[…]
(d) Composition
Figure 2: Edge topological patterns. We refer to the dashed edge(s) as counterpart edge of $(h,r,t)$.
These edge topological patterns are illustrated in Figure 2. It is important to note that a triple can satisfy multiple topological patterns simultaneously, or none at all.
While the above cardinalities and topological patterns have an intuitive definition for individual triples, previous studies (Jin et al., 2023; Teneva & Hruschka, 2023; Bordes et al., 2013; Sun et al., 2019) have considered them as properties of relation types as a whole (e.g., relation $r$ is said to be symmetric if $\>\forall h,t,\;\displaystyle{(h,r,t)\Rightarrow(t,r,h)}$), despite the fact that a given property might be satisfied only by a fraction of the triples of a certain relation. We find that such conceptual generalization lacks a proper formalization in the literature, with inconsistencies in how it is performed, and is prone to introducing noise in results (see Section 5.2). Thus, we consider the above defined properties only as properties of triples.
4 Experimental Setup
We investigate five public biomedical KGs: Hetionet (Himmelstein et al., 2017), PrimeKG [^2] In the original PrimeKG graph, for every triple $(h,r,t)$ the reverse triple $(t,r,h)$ is present as well. We prepocessed PrimeKG to remove these reverse edges. (Chandak et al., 2023), PharmKG (Zheng et al., 2021), OpenBioLink (Breit et al., 2020), PharMeBINet (Königs et al., 2022) (detailed in Table A.1). We also include the trivia KG FB15k-237 (Toutanova & Chen, 2015) as a baseline, to detect any results unique to the biomedical domain. On all KGs we train four of the most popular KGE models: TransE (Bordes et al., 2013), DistMult (Yang et al., 2015), RotatE (Sun et al., 2019) and TripleRE (Yu et al., 2022). Whilst being all well-established among practitioners (Hu et al., 2020), these models differ in complexity and in the relational patterns they can capture (Table A.2). In particular, TransE and DistMult model the interaction between head entity, relation type and tail entity in quite simple terms. Nevertheless, despite missing certain topological patterns, they remain strong baselines. On the other hand, RotatE and TripleRE are more recent approaches able to capture all four investigated topological patterns. Notice however that the theoretical capability of a scoring function to model a particular edge topological pattern is not per se a guarantee of stronger predictive performance on such edges (Jin et al., 2023). Our experiments are designed to quantify this impact.
The training scheme and hyperparameter optimisation details are presented in Appendix B. The results reported in the following sections refer to tail predictions generated on the held out test split, by scoring each $(h,r,?)$ query against all entities in the KG and computing the rank of the ground truth tail $t$, after masking out scores of other $(h,r,t^{\prime})$ triples contained in the graph.
5 Results
5.1 Topological Properties of Biomedical KGs
We first investigate how prevalent the different topological patterns, as described in Section 3, are in the range of public biomedical KGs we consider. Analysing this, we find that many-to-many is the predominant edge cardinality in all considered KGs, with only scarce occurrence of one-to-one, one-to-many, and many-to-one triples (Figure 3). This is in line with the majority of triples having an head out-degree and tail in-degree larger than one (Figures 4 and A.1).
Figure 3: Occurrence of edge cardinalities in the datasets.
Figure 4: Relative frequency of triples when grouped by head out-degree and tail in-degree of the same relation type.
Furthermore, we observe that edge topological patters vary in frequency across the different KGs (Table 1). While composition is the most prevalent pattern in all examined KGs, other edge topological patters occur at a high frequency in only a subset of KGs. Of note is the lack of inverse edges in all the biomedical KGs, apart from OpenBioLink, highlighting that biomedical knowledge is typically encoded in a directed manner in these graphs
Table 1: Occurrence of edge topological patterns as fraction of total triples in the datasets. Note that, since PrimeKG was pre-processed to remove reverse edges as described in Section 4, it does not contain any symmetric triples.
Graph Sym Inf Inv Comp
Hetionet $0.002$ $0.124$ $0.001$ $0.693$
OpenBioLink $0.317$ $0.372$ $0.359$ $0.840$
[…]
PharmKG $0.197$ $0.124$ $0.059$ $0.651$
[…]
FB15k-237 $0.113$ $0.161$ $0.217$ $0.645$
5.2 Effect of Topological Properties on Model Accuracy
Overall, we observe a significant variance in mean reciprocal rank (MRR) achieved between the different KGs as well as between KGE models (Figure 5). In an attempt to understand the root cause of this variance, we focus our analysis on the differences in edge cardinalities and topological patterns described in Section 5.1. When only considering the average occurrence of these topological properties per dataset, the data supports no conclusion about the effect on model accuracy of either edge cardinalities (Figure 1(a)) or edge topological patterns (Figure 1(b)). Previous work (e.g., Teneva & Hruschka (2023)) went one step further, classifying relation types based on the predominant edge topological pattern and cardinality and evaluating the model accuracy per relation type. However, this also does not result in a clear connection between topological properties and KGE model accuracy (Figure C.2). We hypothesize that this could be due the confounding effects of covariates, such as node degrees and different topological patterns, that are too difficult to disentangle, as these properties are often not homogenous within a given relation type. Therefore, unlike prior work, we dissect the link between topological properties and model accuracy at the level of individual triples to allow for a finer-grained analysis with an improved statistical power and a reduced impact of covariates.
Figure 5: Mean reciprocal rank on the test split achieved by different KGE models, for the six datasets.
When comparing the edge cardinality of individual triples to the MRR achieved by the models on these triples, the observed effect strongly differs between the investigated KGs (Figure 6). This result suggests that the exact entity degrees, together with other topological properties, might be better suited to explain the MRR than a binary one/many cardinality classification.
Figure 6: Effect of edge cardinality on MRR.
Indeed, for a given triple, we observe a strong positive correlation between model accuracy and the in-degree of the tail entity as well as a strong negative correlation between model accuracy and the out-degree of the head entity (Figure C.3). This is observed even more clearly when considering the degrees of same relation type (Figures 7 and C.4; see Figures 4 and A.1 for the distribution of degrees in the different KGs). In fact, a high in-degree of the tail node in a tail prediction task has been linked to a higher score (Bonner et al., 2022b), therefore increasing the likelihood of the model predicting it, as confirmed in Figure C.5. On the other hand, a high out-degree of same relation of the head node in a query $(h,r,?)$, implies multiple correct tail entities. Due to the incompleteness of the KG, some of these correct triples might not be present in the graph and thus cannot be filtered out during inference, making the task of predicting the specific entity that is expected by the model more challenging.
Figure 7: Effect of the head and tail degrees of same relation type on MRR.
Consequently, in an attempt to adjust for entity degree, we investigate the effect of edge topological patterns within sets of triples binned by head and tail degrees. Overall, we find that the impact of edge topological patterns on model performance is more relevant if the degrees of head and tail entities are small. When this is the case, compositions are beneficial to MRR across all datasets and models (Figure 8 and Figure C.6). Interestingly, this is true even for DistMult that can’t explicitly model compositional patterns.
Figure 8: The effect of having compositions on MRR, triples grouped by their head and tail degrees of same relation.
When analysing the effect of the patterns is symmetric, has inference and has inverse, we need to distinguish whether the counterpart of the tested edge was present in the training split. If so, the prediction typically becomes much easier due to the availability of additional information seen during training (Toutanova & Chen, 2015), as shown in Figure C.7. An exception to this rule is the prediction of symmetric triples with the scoring function TransE: since TransE by design cannot model symmetric relations (Table A.2), it performs poorly on symmetric triples independently of the counterpart being present in the training data or not. This factor is likely contributing to the subpar accuracy of TransE on OpenBioLink and PharmKG, which have a large fraction of symmetric triples (Figures 5 and 1). On the other hand, our experiments confirm an above average performance of DistMult (which models all relations as symmetric) on symmetric triples, see Figure C.7. In the more challenging case, where the counterpart edge is not present in the training data, symmetry shows a detrimental effect on MRR whereas having inverse or inference has only little impact on accuracy (Figure C.7).
Note that all the above remarks apply equally to biomedical KGs and the general-domain KG FB15k-237.
5.3 Predicting Specific Relation Types
To apply the analysis conducted in the previous sections to real-world tasks, we focus on the relations between pairs of entity types that are of particular interest to practitioners (Gene-Gene, Drug-Gene and Drug-Disease). We investigate how they are represented in different biomedical KGs, and how these differences affect the predictive performance of KGE models. Statistics for the interactions of the considered entity types are given in Figure C.8, while MRR is plotted in Figure 9.
One basic way of characterising the performance of a KGE model is to consider if it can rank entities of the correct type above those of other types when predicting a specific relation (for instance, gene entities when predicting drug-gene interactions) – a task we refer to as demixing. As displayed in Figure C.9, the trained models tend to consistently predict entities of the correct type, albeit with different levels of accuracy. Interestingly, DistMult often shows worse demixing capabilities than other scoring functions. As a consequence of the generally strong demixing performance, the size of the potential set of tails for a given interaction needs to be taken into consideration when comparing MRR across different datasets and relation types (the smaller the set of candidates, the higher the expected MRR achieved by random ranking).
Gene-Gene. As shown in Figure 8(a), these interactions are characterized by a large number of potential tails and a large head out-degree of same relation (with the exception of PharmKG for the latter), making predictions hard. However, a high MRR is observed for PharmKG and OpenBioLink: this is explained by the fact that, for these datasets only (Figure 8(b)), most triples are symmetric (and a significant portion also have inverse/inference edges), with the counterpart likely to have been seen during training (Section 5.2). This also explains the relative ordering of scoring functions for OpenBioLink and PharmKG, with DistMult being the best and TransE the worst.
Drug-Gene. In Hetionet, PrimeKG and OpenBioLink these, interactions are easier to predict for the models than gene-gene interactions (Figure 9). This is likely due to an overall smaller number of potential tails seen during training and, in the case of PrimeKG, a markedly smaller head out-degree. PharmKG and PharMeBINet do not satisfy either of these properties, and indeed we observe no improvements in MRR. The strong predictive power observed for OpenBioLink can be linked to the presence of inverse edges for the vast majority of these triples.
Drug-Disease. Despite a small number of training triples (which is reflected in the sub-optimal demixing profile, Figure 9(c)), drug-disease interactions are easily predicted in Hetionet as only 91 disease entities appear as tails of such triples. Remarkably good predictive performance across all scoring functions is observed for PrimeKG, where the number of candidate tails is larger but the out-degree of head entities remains contained. When both these parameters increase (as in PharMeBINet and PharmKG) performance visibly degrades, despite larger in-degree of tails.
Figure 9: MRR of specific interaction types.
5.4 Case Study: Effect of Additional Training Data
Even when focussing on specific interactions that are represented in multiple KGs, a direct comparison of KGE predictive power across different datasets is limited by having to compare different test sets. In the case of Hetionet and PharMeBINet, however, the latter is constructed by augmenting the former with data coming from other biological databases (Königs et al., 2022). One can therefore often connect individual triples in Hetionet to their exact match in the larger PharMeBINet, which opens up the possibility of studying at a more fundamental level the impact of additional training data on the predictive performance of KGE models. This is important to practitioners as they construct the training KG for a specific task, which is usually done by sub-sampling relevant triples from larger (often proprietary) databases (Chandak et al., 2023).
We consider eight relation types where we can find a significant number of common triples between the two datasets (statistics are given in Table C.1). We extract 10% of the shared triples of the considered relation type as test set and use all other edges in each graph for training. In the case of Hetionet, in addition to the embedding size selected in Section 5.2 maximizing memory utilization, we also train models with the same embedding size chosen for PharMeBINet, which is generally strictly smaller (Hetionet_max and Hetionet_same in Figure 10; hyperparameters in Table B.1). To further ensure a fair comparison, at test time we restrict predictions to a custom set of candidate tails that is the same for the two datasets, namely the set of entities appearing as tails in shared triples of the given relation type.
Figure 10: Comparison of models trained on Hetionet and PharMeBINet when testing MRR on a set of common edges.
As shown in Figures 10 and C.10, across all tested relation types there is no indication that the KGE models are able to benefit from the additional data seen when training on the larger PharMeBINet graph. On the contrary, models trained on Hetionet consistently perform better, even when using the same embedding size. Interestingly, the gap in MRR varies strongly across relation types and also across scoring functions, with DistMult showing generally little difference when comparing models trained with the same embedding size, while distance-based scoring functions experience worse degradation. From Table C.1 we notice that the relation types where the MRR gap is more significant tend to have a larger overall head out-degree (and more relation types coming out from head nodes) for triples in PharMeBINet, compared to Hetionet. We hypothesise that the fact that these nodes are used as head entities for more relation types and triples, many of which are likely not relevant for the specific task at hand, negatively impacts the quality of embeddings. On the other hand, Compound-downregulates-Gene and Compound-causes-Side Effect exemplify relation types where PharMeBINet contains far more triples in addition to the ones in Hetionet. Even in this case, where we could expect the additional training data in the larger dataset to be strictly relevant to the prediction task, all scoring functions show markedly degraded performance when trained on PharMeBINet.
We note that the intention of these results is not to state that one dataset is superior to the other, but rather to highlight any observed differences in how models are able to use training data. However, these results do suggest that, in scenarios where the memory budget is fixed, training on smaller, tailored graphs and increasing the embedding size could be more beneficial than expanding the size of the KG, as the additional data can be a source of confusion for shallow KGE models.
6 Conclusions
This paper analyses the topological properties of widely used biomedical KGs and compares the corresponding predictive performance of different KGE models, focussing in particular on link-prediction tasks relevant to practitioners. Deviating from previous studies, we justify the need to look at properties of individual edges when trying to explain results, as pooling at the relation level introduces too much noise to detect any patterns. By going beyond the coarse one/many binary classification typically used for edge cardinality, we find that considering the actual degrees of head and tail nodes gives a stronger predictor explaining model accuracy. Interestingly, similar interpretations apply to both biomedical KGs and the general-domain KG. Edge topological patterns also impact predictive accuracy, especially when entity degrees are small. For such patterns, we observe an improved accuracy when the counterpart edge (e.g., the reverse edge for symmetric triples) has been seen during training. This in turn raises the problem of inconsistent model validation schemes and topological property definitions when comparing results of previous studies on similar datasets. We address this problem by releasing all predictions from our experiments, together with a new toolbox for KG topological analysis. Finally, by performing a case study comparing predictions on common sets of edges shared by different KGs, we show that training on larger graphs, encoding more biomedical data, can unintuitively harm predictive performance. This evidence should encourage a wider discussion on the guiding principles to adopt when constructing the training KG for biomedical tasks – a crucial problem for real-world practitioners, but scarcely investigated in the literature.
Appendix A Dataset and Model Properties
Table A.1: Properties of the considered datasets.
Graph # Entities # Relations # Triples Avg node degree
Hetionet 45,158 24 2,250,197 99.66
OpenBioLink 184,635 28 4,563,405 49.43
PharMeBINet 2,653,751 208 15,883,653 11.97
PharmKG 188,296 39 1,093,236 11.61
PrimeKG 129,375 30 4,050,064 62.61
FB15k-237 14,541 237 310,116 42.65
Figure A.1: Relative frequency of triples when grouped by head out-degree and tail in-degree of the same relation type.
Table A.2: Scoring functions and their ability to model four fundamental relation properties: S = Symmetry; INF = Inference; INV = Inversion; C = Composition. For RotatE we assume $d$ even and denote by $\mathbb{C}^{\frac{d}{2}}$ the vector space $\mathbb{R}^{d}=(\mathbb{R}\oplus i\mathbb{R})^{\frac{d}{2}}$ with the structure of $\mathbb{R}$-algebra induced by the product of complex numbers. $\circ$ denotes the Hadamard product; $p\in\left\{1,2\right\}$.
Model Scoring function S INF INV C
TransE $-\left\lVert{\bm{h}}+{\bm{r}}-{\bm{t}}\right\rVert_{p}$ ${\bm{h}},{\bm{r}},{\bm{t}}\in\mathbb{R}^{d}$
RotatE $-\left\lVert{\bm{h}}\circ e^{i{\bm{r}}}-{\bm{t}}\right\rVert_{p}$ ${\bm{h}},{\bm{t}}\in\mathbb{C}^{\frac{d}{2}},{\bm{r}}\in\mathbb{R}^{\frac{d}{2}}$
DistMult $\langle{\bm{r}},{\bm{h}},{\bm{t}}\rangle$ ${\bm{h}},{\bm{r}},{\bm{t}}\in\mathbb{R}^{d}$
TripleRE $-\left\lVert{\bm{h}}\circ\bm{r_{h}}+\bm{r_{m}}-{\bm{t}}\circ\bm{r_{t}}\right\rVert$ ${\bm{h}},\bm{r_{h}},\bm{r_{m}},{\bm{t}},\bm{r_{t}}\in\mathbb{R}^{d}$
Appendix B Details on Experimental Setup and Hyperparameter Selection
[…]
Table B.1: Experiment hyperparameters for different datasets. Hetionet_same refers to the alternative experimental configuration used in Section 5.4.
Graph Model Hidden size Learning Rate Scoring Norm # Negative samples / positive
Hetionet DistMult $2048$ $0.0003$ $16$
[…]
Hetionet_same DistMult $300$ $0.0003$ $16$
[…]
OpenBioLink DistMult $768$ $0.0003$ $16$
[…]
PharMeBINet DistMult $300$ $0.003$ $16$
[…]
PharmKG DistMult $768$ $0.003$ $16$
[…]
PrimeKG DistMult $1024$ $0.0003$ $16$
[…]
FB15k-237 DistMult $4096$ $0.001$ $16$
[…]
Appendix C Additional Results
Figure C.1: Mean reciprocal rank plotted against the average occurrence of edge cardinalities (a) and edge topological patterns (b) in the six datasets.
Figure C.2: Spearman-rank correlation between the average MRR of a relation type and the average frequency of edge cardinalities (a) and of topological patterns (b) in that relation type.
Figure C.3: Spearman-rank correlation between MRR of individual triples and the out-degree of the head node as well as in-degree of the tail node.
(a) PharMeBINet
(b) PharmKG
(c) PrimeKG
(d) FB15k-237
Figure C.4: Effect of head out-degree and tail in-degree on MRR for additional datasets.
Figure C.5: Distribution of Spearman-rank correlation between in-degree of same relation and how frequently the entity is incorrectly selected among the top-$100$ tail predictions, grouping test queries by relation type. A positive correlation means that KGE models are biased towards predicting entities with a larger number of incoming edges of the relation type considered in the query.
(a) PharMeBINet
(b) PharmKG
(c) PrimeKG
(d) FB15k-237
Figure C.6: The effect of having compositions on MRR, triples grouped by their head and tail degrees.
Figure C.7: The effect on MRR of being symmetric and having inference/inverse, distinguishing based on whether the counterpart edge is present or absent in in the training data. Note that symmetric triples are too rare in PharMeBINet for a meaningful analysis. Triples grouped by their head and tail degrees.
(a) Median head out-degree and tail in-degree of same relation and number of unique relation tail entities.
(b) Frequency of edge patterns.
Figure C.8: Statistics of topological properties for the interaction types investigated in Section 5.3.
(a) Gene-Gene
(b) Drug-Gene
(c) Drug-Disease
Figure C.9: Demixing for the interaction types investigated in Section 5.3. For each test query, we compute how many of the top-$100$ predictions made by the model are contained in the set of entities used as tails by triples of the considered relations.
Table C.1: Statistics of the relation types compared between Hetionet and PharMeBINet in Section 5.4. Matching rate denotes the fraction of triples of the relation type that are matched to triples in the other dataset. For the rows Unique heads/tails/out-relations/in-relations and Head/Tail out-/in-degree (same relation), the median value is reported. For the rows Has inverse/inference/composition, we report the fraction of total relation triples with the given property.
Relation Disease-localizes-Anatomy Compound-binds-Gene Gene-covaries-Gene Anatomy-expresses-Gene
Dataset Hetionet PhMBNet Hetionet PhMBNet Hetionet PhMBNet Hetionet PhMBNet
Relation triples $3602$ $3602$ $11571$ $11622$ $61690$ $61615$ $526407$ $526180$
Matching rate $1.0$ $1.0$ $0.998$ $0.993$ $0.999$ $1.0$ $0.999$ $1.0$
Test triples $360$ $360$ $1141$ $1141$ $6161$ $6161$ $52618$ $52618$
Unique heads $133$ $133$ $1389$ $1426$ $9043$ $9034$ $241$ $241$
Unique tails $398$ $398$ $1689$ $1701$ $9542$ $9518$ $18094$ $18074$
Head out-degree $212$ $227$ $132$ $2085$ $74$ $128$ $11952$ $11945$
Head out-degree s.r. $34$ $34$ $14$ $14$ $20$ $20$ $7937$ $7935$
Unique out-relations $4$ $6$ $4$ $20$ $4$ $8$ $3$ $3$
Tail in-degree $11$ $11$ $102$ $162$ $83$ $114$ $77$ $112$
Tail in-degree s.r. $11$ $11$ $36$ $36$ $17$ $17$ $44$ $44$
Unique in-relations $1$ $1$ $7$ $13$ $6$ $11$ $6$ $11$
Has inverse $0.0$ $0.0$ $0.0$ $0.03$ $0.001$ $0.001$ $0.0$ $0.0$
Has inference $0.0$ $0.0$ $0.006$ $0.08$ $0.002$ $0.002$ $0.263$ $0.263$
Has composition $0.591$ $0.594$ $0.571$ $0.957$ $0.501$ $0.507$ $0.907$ $0.907$
Relation Compound-causes-Side Effect Gene-regulates-Gene Gene-interacts-Gene Compound-downregulates-Gene
Dataset Hetionet PhMBNet Hetionet PhMBNet Hetionet PhMBNet Hetionet PhMBNet
Relation triples $138944$ $154511$ $265672$ $265667$ $147164$ $147133$ $21102$ $231156$
Matching rate $0.909$ $0.817$ $0.999$ $1.0$ $0.999$ $1.0$ $0.997$ $0.098$
Test triples $12630$ $12630$ $26566$ $26566$ $14713$ $14713$ $2105$ $2105$
Unique heads $1071$ $1358$ $4634$ $4634$ $9526$ $9525$ $734$ $2631$
Unique tails $5701$ $6023$ $7048$ $7047$ $14084$ $14073$ $2880$ $21912$
Head out-degree $245$ $2186$ $203$ $254$ $214$ $267$ $515$ $4144$
Head out-degree s.r. $201$ $182$ $104$ $104$ $54$ $54$ $225$ $1413$
Unique out-relations $5$ $19$ $6$ $10$ $5$ $10$ $5$ $20$
Tail in-degree $164$ $254$ $309$ $370$ $106$ $145$ $252$ $116$
Tail in-degree s.r. $164$ $193$ $208$ $208$ $27$ $27$ $20$ $15$
Unique in-relations $1$ $3$ $8$ $12$ $7$ $11$ $8$ $11$
Has inverse $0.0$ $0.0$ $0.003$ $0.003$ $0.006$ $0.006$ $0.0$ $0.002$
Has inference $0.0$ $0.142$ $0.006$ $0.003$ $0.006$ $0.006$ $0.001$ $0.164$
Has composition $0.366$ $0.893$ $0.881$ $0.886$ $0.703$ $0.713$ $0.913$ $0.800$
Figure C.10: MRR comparison for additional relation types, when testing on a set of common edges between Hetionet and PharMeBINet.